OpenStudy (imsleeziboii):

If the slope of a straight line is 0 and the y-intercept is –2, what is the equation of the line?

OpenStudy (imsleeziboii):

HELP!!

OpenStudy (imsleeziboii):

@rebeccaxhawaii

OpenStudy (theprofessor35):

y = mx + b is the form

OpenStudy (theprofessor35):

Lets see. m is the slope, and b is y intercept , so just replace.

OpenStudy (theprofessor35):

If the slope of a straight line is 0 and the y-intercept is –2, what is the equation of the line? y = mx + b y = 0x - 2 y = -2

OpenStudy (4everaddicted2anime):

Everything times 0 is 0, so all you should be left with after plugging everything in and simplifying is y=-2

OpenStudy (theprofessor35):

Since there is no 'x' value here, is basically just the y inetrcept. No slope y = -2

OpenStudy (theprofessor35):

There you go, you have an answer.

OpenStudy (imsleeziboii):

can you medal 4ever for me? i can only medal one but you both helped

OpenStudy (imsleeziboii):

i mean can you medal theproffessor

OpenStudy (theprofessor35):

You can undo a medal I assume.

OpenStudy (4everaddicted2anime):

I did

OpenStudy (imsleeziboii):

But you both helped.